OpenStudy (anonymous):

(4x - 6y = 3, 6x - 4y = -3)

OpenStudy (anonymous):

\[y = \frac{ 2 }{ 3 }x - \frac{ 1 }{ 2 }, y= \frac{ 3}{ 2}x + \frac{ 3 }{ 4 } \]